OpenStudy (anonymous):

please help, im trying to finish a timed quiz, Determine the solution set of (x - 4)^2 = 12. {4 + 2√3, 4 - 2√3} {4 - 2√3, 4 - 2√3} {2√3 + 4, 2√3 - 4}

OpenStudy (anonymous):

you may use square root property on both sides.

OpenStudy (alexwee123):

square root both sides...

OpenStudy (alexwee123):

and honestly if this is a quiz then you should really do it yourself....

OpenStudy (anonymous):

not exactly a quiz, but makeup work for school, online, but it is timed,

OpenStudy (anonymous):

example: \[(x-5)^{2}=\sqrt{18} ===> x-5 = +/- \sqrt{18} ==> x-5 = +/- \sqrt{9}*\sqrt{2} ==> x- 5 = +/- 3\]

OpenStudy (alexwee123):

nonetheless it's still an assignment and this is cheating somewhat :/

OpenStudy (anonymous):

we can give her an example instead. :))

OpenStudy (anonymous):

continuation of the example: ==>x−5=+/−3

OpenStudy (anonymous):

honestly i dont have time for this, i have so much work it's not even funny, and im have only 3 more days left to finish i dont need examples i need you to help me so i can move on to the next lesson